### v1

@@ Abstract
Code is available at https://github.com/aaa493/DCHO.
keywords:
Higher-Order Brain Connectivity, Brain Dynamic Prediction
@@ 1 Introduction
Functional connectivity (FC) [1, 2], the most widely used framework, models pairwise statistical dependencies between brain regions based on fMRI data. FC is limited to pairwise interactions and may overlook higher-level coordination. Recent studies highlight the importance of higher-order brain connectivity (HOBC) that involves three or more regions, supported by evidence across scales [3, 4, 5, 6, 7, 8, 9, 10]. In human neuroimaging, direct observation of HOBC remains challenging due to the limitations of noninvasive techniques. A recent study [11, 12] was the first to infer latent HOBC from fMRI signals and applied them to downstream tasks such as classification. However, they focus on analyzing HOBC within static or predefined temporal windows, without modeling their temporal evolution. This limits the potential of HOBC in predictive applications, such as forecasting brain dynamics. Addressing this gap requires models that can anticipate the dynamic trajectory of the HOBC over time.
@@
Predicting the evolution of the HOBC in complex brain systems is challenging. First, the combinatorial explosion in the number of possible higher-order interactions poses a major obstacle to direct modeling [13, 14, 15, 16]. The curse of dimensionality poses a major challenge to inferring the HOBC. Second, brain dynamics are inherently nonlinear, nonstationary, and stochastic, governed by latent neural processes that evolve across multiple spatial and temporal scales. Capturing the temporal evolution of HOBC thus requires not only modeling individual node trajectories, but also learning coordinated multi-region dependencies that reflect complex integration, segregation, and recurrent feedback mechanisms. These interactions are not additive but shaped by nonlinear dynamical couplings that defy simple compositional rules [17, 18, 19, 20].
@@ 2 Related Work
Higher-order Brain Connectivity: Although FC [1, 2] has been the standard for brain network modeling, it fails to capture higher-order interactions among multiple brain regions. HOBC describes complex coordination among three or more regions and reveals richer organizational patterns. However, due to the limitations of non-invasive neuroimaging techniques, directly observing HOBC in humans remains challenging. Recent work [11, 12] has attempted to infer HOBC from fMRI and apply it to static classification tasks. However, the study overlooked its temporal dynamics, limiting their use in predictive settings. To bridge this gap, we propose the DCHO, which models and forecasts the temporal evolution of HOBC from neural signals, enabling both non-predictive tasks (task state classification) and more demanding predictive tasks (brain dynamics forecasting).
Brain Dynamics Forecasting: In recent years, various architectures have been proposed for modeling neural dynamics. RNN-based methods (LtrRNN [21] and LSTM [22]) are effective in capturing short-term temporal dependencies but are limited in expressing long-term dynamics. Transformer-based models (NetFormer [23], STNDT [24] and Transformer [25]) leverage self-attention mechanisms to better model global temporal dependencies, yet often overlook the underlying brain connectivity structure. GNN-based methods (AMAG [26]) are naturally suited for encoding spatial topological structures between brain regions and have been applied to model static or pairwise functional connectivity, but typically neglect global topological information. In contrast, the proposed DCHO framework integrates a dual-graph encoder and a latent learner to capture high-level information representations of HOBC and explicitly models the underlying dynamic evolution in the latent space, substantially enhancing modeling capacity and robustness in long-term prediction.
@@ 3.1 The Definition of Higher-order Brain Connectivity
We first introduce the definition of HOBC [11, 12]. Let $\mathbf{x}_{i}\in\mathbb{R}^{T}=[x_{i}^{1},x_{i}^{2},\ldots,x_{i}^{T}]$ denotes the original time series of region $i$. We first $z$-score each region’s time series:
@@ 4.2 The Dual-view Encoder
(i) The local topological extractor adaptively encodes pairwise interactions using a GNN-based mechanism [27, 28]. First, interaction scores are computed between the central node and its neighbors, and these scores are used to weight the neighbors’ embeddings from the previous layer. Mathematically, for the embedding of the $i_{t}$ node $v_{i}^{t}$ at the $s$-th layer: $e_{i}^{t,(s)}$, the interaction scores between the $i_{t}$ node and its neighbor $j_{t}$ are derived from the adjacency matrix and the node features in the temporal graph:
@@
(ii) The global topological extractor is designed based on spectral graph convolution, which serves as a complementary approach to uncover HOBC information hidden in the spectral domain. Specifically, first, we introduce the necessary mathematical definitions. The Chebyshev polynomials are defined as $T_{0}(x)=1$, $T_{1}(x)=x$, and $T_{m}(x)=2xT_{m-1}(x)-T_{m-2}(x)$. The normalized graph Laplacian $L$ is defined as $L=I-\widetilde{D}^{-\frac{1}{2}}A\widetilde{D}^{-\frac{1}{2}}$, where $D$ is the degree matrix and $I$ is the identity matrix. Define the matrix $R^{(0)}=X$, which stacks the features of all nodes. Then, based on the second-order Chebyshev graph convolution [29], the node representation matrix $R^{(s)}$ at the $s$-th layer is calculated as:
@@
(iii) The above two extractors have captured the representations of pairwise interactions and higher-order interactions, respectively. We then employ an attention mechanism to aggregate these temporal representations into a latent representation $\mathbf{Z}^{t-T,t}$ with advanced spatio-temporal features for each object. Specifically, first, we merge the two representations from the last layer with the temporal embeddings $TE(t)$, and apply a MLP $\delta(\cdot)$ to calculate the final representation $c_{i}^{t}$ for each node $v_{i}^{t}$. Mathematically:
@@
Then we summarize the representations of the nodes over all timestamps into a compact representation $\mathbf{Z}_{i}^{t-T,t}$ by generating weights for each timestamp and calculating the weighted average. $\mathbf{Z}_{i}^{t-T,t}$ represents the latent representation that encapsulates high-level spatio-temporal information. Mathematically:
$\zeta_{i}^{n}$ $=\tanh\left(\left(\frac{1}{T}\sum_{n=t-T}^{t}c_{i}^{n}\right)W_{\text{sum}}\cdot c_{i}^{n}\right),$ (15)
$\mathbf{Z}_{i}^{t-T,t}$ $=\frac{1}{T}\sum_{n=t-T}^{t}\sigma\left(\zeta_{i}^{n}c_{i}^{n}\right),$ (16)
where $W_{\text{sum}}$ is a trainable matrix designed for linear transformation, and $\sigma(\cdot)$ refers to a non-linear activation function.
@@ 4.4 LSTM-based predictor
5 Results
5.1 Dataset and Preprocessing
To evaluate the performance of our proposed method, we conducted experiments on both synthetic data and real-world fMRI datasets. The synthetic data were generated from two nonlinear dynamical systems: the Lorenz system [30] and the Hindmarsh–Rose (HR) neuronal model [31]. In both cases, we simulated networks consisting of 20 coupled nodes. For real-world evaluation, we utilized the Human Connectome Project (HCP) dataset [32], which includes both resting-state and task-based fMRI recordings. Specifically, we employed seven task-based datasets: Emotion, Gambling, Language, Motor, Relational, Social, and Working Memory (WM) and a resting-state dataset (Rest). All HCP data had been preprocessed using the standard minimal preprocessing pipeline. We parcellated the cortex into 17 regions based on the Yeo 17-network atlas. It is important to note that the dimensionality of the higher-order brain connectivity tensor $H^{t}\in\mathbb{R}^{N\times N\times N}$ increases cubically with the number of brain regions N, indicating that the HOBC derived from the Yeo-17 parcellation is already highly complex and high-dimensional. For each subject and session, we extracted the average BOLD signal within each region and applied temporal standardization (z-scoring) independently for each region. In terms of data segmentation, all time series were divided into fixed-length temporal segment using a sliding window, with the window length set equal to the prediction length and a stride of 3. Following segmentation, 80% of the data was used for training and the remaining 20% for testing. See Appendix C and Appendix E for further details.
Experimental Setup: Since HOBC prediction is a novel task, no existing methods have been specifically designed for this challenge. We therefore compare DCHO with commonly used baselines, including MLP [33], LSTM [22], and Transformer [25], to provide a meaningful reference. Detailed descriptions of these methods are provided in Appendix D. We evaluate model performance using two common metrics (MAE and RMSE) across ten datasets (Emotion, Gambling, Language, Motor, Relational, Social, WM, Rest, Lorenz, HR).
@@ 5.4 Predictive Downstream Task
Compared to the study [11, 12] that focus primarily on analyzing HOBC within static or predefined temporal windows, DCHO captures the underlying dynamic evolution of HOBC, enabling more sophisticated predictive tasks.
Benchmark Methods and Evaluation Metrics: To evaluate the effectiveness of the pretrained dual-graph encoder and predictor, we compare DCHO against several state-of-the-art methods, including RNN-based models (LtrRNN [21], LSTM [22]), Transformer-based models (Transformer [25], STNDT [24] and NetFormer [23]), GNN-based models (AMAG [26]), and a standard MLP [33] baseline. Detailed descriptions of these methods are provided in Appendix D. We evaluate model performance using two common metrics (MAE and RMSE) across ten datasets (Emotion, Gambling, Language, Motor, Relational, Social, WM, Rest, Lorenz, HR).
@@ Appendix D Details of baseline
  • MLP [33]: A standard multi-layer perceptron that models temporal dynamics, without considering spatial or temporal structure explicitly.
  • LSTM [22]: A recurrent neural network that captures temporal dependencies via gating mechanisms, widely used in sequence modeling tasks.
  • Transformer [25]: A self-attention-based model that captures long-range dependencies in time series, without relying on recurrence.
  • STNDT [24]: A Transformer-based method, STNDT captures spatiotemporal neural dynamics and enhances prediction through contrastive learning.
  • AMAG [26]: A graph-based method, AMAG uses additive and multiplicative message passing to adaptively model pairwise interactions and forecast neural activity.
  • LtrRNN [21]: A tensor-based model that uncovers low-dimensional neural dynamics by fitting RNNs to large-scale cortical recordings during learning tasks.
  • NetFormer [23]: A transformer-based model designed for interpretable prediction of nonstationary neuronal dynamics and time-varying connectivity patterns from population activity.

### v2

@@ Abstract
Codehttps://github.com/aaa493/DCHO
Datahttps://huggingface.co/datasets/aaa493/DCHO
@@ 1 Introduction
Functional connectivity (FC) (yeo2011organization; hutchison2013dynamic), the most widely used framework, models pairwise statistical dependencies between brain regions based on fMRI data. FC is limited to pairwise interactions and may overlook higher-level coordination. Recent studies highlight the importance of higher-order brain connectivity (HOBC) that involves three or more regions, supported by evidence across scales (BATTISTON20201; battiston2022higher; bianconi2021higher; jun2017fully; steinmetz2021neuropixels; paulk2022large; chelaru2021high; tadic2022multiscale). In human neuroimaging, direct observation of HOBC remains challenging due to the limitations of noninvasive techniques. A recent study (santoro2024higher; santoro2023higher) was the first to infer latent HOBC from fMRI signals and applied them to downstream tasks such as classification. However, they focus on analyzing HOBC within static or predefined temporal windows, without modeling their temporal evolution. This limits the potential of HOBC in predictive applications, such as forecasting brain dynamics. Addressing this gap requires models that can anticipate the dynamic trajectory of the HOBC over time.
@@
Predicting the evolution of the HOBC in complex brain systems is challenging. First, the combinatorial explosion in the number of possible higher-order interactions poses a major obstacle to direct modeling (bassett2017network; breakspear2017dynamic; ROEBROECK2011296; HERZOG2022105918). The curse of dimensionality poses a major challenge to inferring the HOBC. Second, brain dynamics are inherently nonlinear, nonstationary, and stochastic, governed by latent neural processes that evolve across multiple spatial and temporal scales. Capturing the temporal evolution of HOBC thus requires not only modeling individual node trajectories, but also learning coordinated multi-region dependencies that reflect complex integration, segregation, and recurrent feedback mechanisms. These interactions are not additive but shaped by nonlinear dynamical couplings that defy simple compositional rules (Topologyshapes; battiston2021physics; zhang2023higher; alvarez2021evolutionary).
@@ 2 Related Work
Higher-order Brain Connectivity: Although FC (yeo2011organization; hutchison2013dynamic) has been the standard for brain network modeling, it fails to capture higher-order interactions among multiple brain regions. HOBC describes complex coordination among three or more regions and reveals richer organizational patterns. However, due to the limitations of non-invasive neuroimaging techniques, directly observing HOBC in humans remains challenging. Recent work (santoro2024higher; santoro2023higher) has attempted to infer HOBC from fMRI and apply it to static classification tasks. However, the study overlooked its temporal dynamics, limiting their use in predictive settings. To bridge this gap, we propose the DCHO, which models and forecasts the temporal evolution of HOBC from neural signals, enabling both non-predictive tasks (task state classification) and more demanding predictive tasks (brain dynamics forecasting).
Brain Dynamics Forecasting: In recent years, various architectures have been proposed for modeling neural dynamics. RNN-based methods (LtrRNN (pellegrino2023low) and LSTM (graves2012long)) are effective in capturing short-term temporal dependencies but are limited in expressing long-term dynamics. Transformer-based models (NetFormer (lu2025netformer), STNDT (le2022stndt) and Transformer (vaswani2017attention)) leverage self-attention mechanisms to better model global temporal dependencies, yet often overlook the underlying brain connectivity structure. GNN-based methods (AMAG (li2023amag)) are naturally suited for encoding spatial topological structures between brain regions and have been applied to model static or pairwise functional connectivity, but typically neglect global topological information. In contrast, the proposed DCHO framework integrates a dual-graph encoder and a latent learner to capture high-level information representations of HOBC and explicitly models the underlying dynamic evolution in the latent space, substantially enhancing modeling capacity and robustness in long-term prediction.
@@ 3.1 The Definition of Higher-order Brain Connectivity
We first introduce the definition of HOBC (santoro2024higher; santoro2023higher). Let $\mathbf{x}_{i}\in\mathbb{R}^{T}=[x_{i}^{1},x_{i}^{2},\ldots,x_{i}^{T}]$ denotes the original time series of region $i$. We first $z$-score each region’s time series:
@@ 4.2 The Dual-view Encoder
(i) The local topological extractor adaptively encodes pairwise interactions using a GNN-based mechanism (kipf2017semisupervisedclassificationgraphconvolutional; xu2019powerful). First, interaction scores are computed between the central node and its neighbors, and these scores are used to weight the neighbors’ embeddings from the previous layer. Mathematically, for the embedding of the $i_{t}$ node $v_{i}^{t}$ at the $s$-th layer: $e_{i}^{t,(s)}$, the interaction scores between the $i_{t}$ node and its neighbor $j_{t}$ are derived from the adjacency matrix and the node features in the temporal graph:
@@
(ii) The global topological extractor is designed based on spectral graph convolution, which serves as a complementary approach to uncover HOBC information hidden in the spectral domain. Specifically, first, we introduce the necessary mathematical definitions. The Chebyshev polynomials are defined as $T_{0}(x)=1$, $T_{1}(x)=x$, and $T_{m}(x)=2xT_{m-1}(x)-T_{m-2}(x)$. The normalized graph Laplacian $L$ is defined as $L=I-\widetilde{D}^{-\frac{1}{2}}A\widetilde{D}^{-\frac{1}{2}}$, where $D$ is the degree matrix and $I$ is the identity matrix. Define the matrix $R^{(0)}=X$, which stacks the features of all nodes. Then, based on the second-order Chebyshev graph convolution (defferrard2016convolutional), the node representation matrix $R^{(s)}$ at the $s$-th layer is calculated as:
@@
(iii) The above two extractors have captured the representations of pairwise interactions and higher-order interactions, respectively. We then employ an attention mechanism to aggregate these temporal representations into a latent representation $\mathbf{Z}_{i}^{t}$ with advanced spatio-temporal features. Specifically, first, we merge the two representations from the last layer with the temporal embeddings $TE(t)$, and apply a MLP $\delta(\cdot)$ to calculate the representation $c_{i}^{t}$ for each node $v_{i}^{t}$. Mathematically:
@@
We then refine the node representation through an attention-based transformation that integrates contextual dependencies at each timestamp. Specifically, an attention refinement operator is applied to produce a context-aware intermediate state, which is further processed by a position-wise feed-forward network to obtain the final latent representation. Mathematically:
$\zeta_{i}^{t}$ $=\psi(c_{i}^{t}+f_{att}(c_{i}^{t})),$ (15)
$\mathbf{Z}_{i}^{t}$ $=\psi(\zeta_{i}^{t}+f_{ffn}(\zeta_{i}^{t})).$ (16)
where $f_{\mathrm{att}}(\cdot)$ denotes an attention refinement operator implemented with multi-head self-attention, $f_{\mathrm{ffn}}(\cdot)$ is a position-wise feed-forward network, and $\psi(\cdot)$ denotes a normalization operator (Layer Normalization ).
@@ 4.4 LSTM-based predictor
5 Results
5.1 Dataset and Preprocessing
To evaluate the performance of our proposed method, we conducted experiments on both synthetic data and real-world fMRI datasets. The synthetic data were generated from two nonlinear dynamical systems: the Lorenz system (wang2020adaptive) and the Hindmarsh–Rose (HR) neuronal model (hindmarsh1984model). In both cases, we simulated networks consisting of 20 coupled nodes. For real-world evaluation, we utilized the Human Connectome Project (HCP) dataset (van2012human), which includes both resting-state and task-based fMRI recordings. Specifically, we employed seven task-based datasets: Emotion, Gambling, Language, Motor, Relational, Social, and Working Memory (WM) and a resting-state dataset (Rest). All HCP data had been preprocessed using the standard minimal preprocessing pipeline. We parcellated the cortex into 17 regions based on the Yeo 17-network atlas. It is important to note that the dimensionality of the higher-order brain connectivity tensor $H^{t}\in\mathbb{R}^{N\times N\times N}$ increases cubically with the number of brain regions N, indicating that the HOBC derived from the Yeo-17 parcellation is already highly complex and high-dimensional. For each subject and session, we extracted the average BOLD signal within each region and applied temporal standardization (z-scoring) independently for each region. In terms of data segmentation, all time series were divided into fixed-length temporal segment using a sliding window, with the window length set equal to the prediction length and a stride of 3. Following segmentation, 80% of the data was used for training and the remaining 20% for testing. See Appendix C and Appendix E for further details.
Experimental Setup: Since HOBC prediction is a novel task, no existing methods have been specifically designed for this challenge. We therefore compare DCHO with commonly used baselines, including MLP (rosenblatt1958perceptron), LSTM (graves2012long), and Transformer (vaswani2017attention), to provide a meaningful reference. Detailed descriptions of these methods are provided in Appendix D. We evaluate model performance using two common metrics (MAE and RMSE) across ten datasets (Emotion, Gambling, Language, Motor, Relational, Social, WM, Rest, Lorenz, HR).
@@ 5.4 Predictive Downstream Task
Compared to the study (santoro2024higher; santoro2023higher) that focus primarily on analyzing HOBC within static or predefined temporal windows, DCHO captures the underlying dynamic evolution of HOBC, enabling more sophisticated predictive tasks.
Benchmark Methods and Evaluation Metrics: To evaluate the effectiveness of the pretrained dual-graph encoder and predictor, we compare DCHO against several state-of-the-art methods, including RNN-based models (LtrRNN (pellegrino2023low), LSTM (graves2012long)), Transformer-based models (Transformer (vaswani2017attention), STNDT (le2022stndt) and NetFormer (lu2025netformer)), GNN-based models (AMAG (li2023amag)), and a standard MLP (rosenblatt1958perceptron) baseline. Detailed descriptions of these methods are provided in Appendix D. We evaluate model performance using two common metrics (MAE and RMSE) across ten datasets (Emotion, Gambling, Language, Motor, Relational, Social, WM, Rest, Lorenz, HR).
@@ 6 Conclusion
Acknowledgments
This work was supported by the National Natural Science Foundation of China (62472206), National Key R&D Program of China (2025YFC3410000), Shenzhen Science and Technology Innovation Committee (RCYX20231211090405003, KJZD20230923115221044), GuangDong Basic and Applied Basic Research Foundation (2025A1515011645 to ZC.L.), Shenzhen Doctoral Startup Project (RCBS20231211090748082 to XK.S.), Guangdong Provincial Key Laboratory of Advanced Biomaterials (2022B1212010003), and the open research fund of the Guangdong Provincial Key Laboratory of Mathematical and Neural Dynamical Systems, the Center for Computational Science and Engineering at Southern University of Science and Technology.
@@ Appendix D Details of baseline
  • MLP (rosenblatt1958perceptron): A standard multi-layer perceptron that models temporal dynamics, without considering spatial or temporal structure explicitly.
  • LSTM (graves2012long): A recurrent neural network that captures temporal dependencies via gating mechanisms, widely used in sequence modeling tasks.
  • Transformer (vaswani2017attention): A self-attention-based model that captures long-range dependencies in time series, without relying on recurrence.
  • STNDT (le2022stndt): A Transformer-based method, STNDT captures spatiotemporal neural dynamics and enhances prediction through contrastive learning.
  • AMAG (li2023amag): A graph-based method, AMAG uses additive and multiplicative message passing to adaptively model pairwise interactions and forecast neural activity.
  • LtrRNN (pellegrino2023low): A tensor-based model that uncovers low-dimensional neural dynamics by fitting RNNs to large-scale cortical recordings during learning tasks.
  • NetFormer (lu2025netformer): A transformer-based model designed for interpretable prediction of nonstationary neuronal dynamics and time-varying connectivity patterns from population activity.
@@ Appendix E Implementation Details
Appendix F Additional ablation experiments
To further validate the effectiveness of our method, we include additional ablation experiments. DCHO.f is a variant that replaces the dual-view encoder with an MLP. DCHO.g substitutes the LSTM with a Transformer, and DCHO.h replaces the LSTM with a GRU. As shown in Table 4, the inferior performance of DCHO.f further confirms the effectiveness of the dual-view encoder. In addition, both DCHO.g and DCHO.h exhibit only slight performance degradation, which further verifies that our model’s gains mainly derive from the decomposition–composition strategy (i.e., the latent prediction loss) rather than from the choice of predictor architecture.
Table 4: Additional ablation analysis
Methods Emotion Rest
DCHO.f 0.3849 $\pm$ 0.0019 0.3527 $\pm$ 0.0013
DCHO.g 0.1841 $\pm$ 0.0007 0.1725 $\pm$ 0.0011
DCHO.h 0.1962 $\pm$ 0.0010 0.1812 $\pm$ 0.0008
DCHO 0.1744 $\pm$ 0.0004 0.1704 $\pm$ 0.0004
Appendix G Expanded Experimental Evaluation with Additional Datasets and Baselines
To further evaluate the performance of our model, we additionally include more baseline methods (HGFM, ALTER, and NEUROTREE) as well as a new dataset (the resting-state subset of CHCP). These baselines were designed based on graph architectures but focus primarily on diagnostic tasks rather than dynamic HOBC prediction, and therefore cannot be directly compared under our task setting. To enable a fair and appropriate comparison, we adapt HGFM, ALTER, and NEUROTREE to our experimental environment. As shown in Table 5, all three baselines exhibit lower performance than our model.
Table 5: Additional experimental evaluation
Methods Emotion (HCP) Rest (CHCP)
HGFM 0.3721 $\pm$ 0.0007 0.2324 $\pm$ 0.0016
ALTER 0.3905 $\pm$ 0.0021 0.2652 $\pm$ 0.0017
NEUROTREE 0.4169 $\pm$ 0.0014 0.2360 $\pm$ 0.0014
DCHO 0.1744 $\pm$ 0.0004 0.2114 $\pm$ 0.0010
Appendix H Complexity analyses
Our method consists of two stages: the first stage infers HOBC, and the second stage performs latent trajectory prediction. Since the second stage operates in a low-dimensional latent space, its computational cost is negligible. Therefore, we focus on the complexity of the first stage. The main cost in the first stage comes from the dual-view encoder and the high-order decoder. The dual-view encoder has two parallel branches: a local topology extractor and a global topology extractor.
Local branch complexity comprises four parts: adjacency normalization: $O(BTN^{2})$, linear projection: $O(BTNF^{2})$, cosine similarity: $O(BTN^{2})$, neighbor aggregation: $O(BTN^{2}F)$. Summing yields
$O\!\left(BT\,(N^{2}F+NF^{2}+N^{2})\right)\approx O(BTNF^{2}),$ (28)
since $NF^{2}>N^{2}F>N^{2}$ in our setting. The global branch includes the normalized Laplacian construction $O(BTN^{2})$, the Chebyshev recurrence $O(BTN^{2}F)$, and the Chebyshev linear mapping $O(BTNF^{2})$. Summing these terms gives
$O\!\left(BT\,(N^{2}+N^{2}F+NF^{2})\right)\approx O(BTNF^{2}).$ (29)
Since the two branches run in parallel, the total complexity is determined by the dominant term among them. Therefore, the overall complexity of the dual-view encoder is $O(BTNF^{2})$.
The high-order decoder consists of five components: linear projection: $O(BTNF^{2})$, latent composition learner: $O(BTMF^{2})$, spatial Transformer: $O(BTM^{2}F)$, temporal Transformer: $O(BTMF^{2})$, output projection: $O(BTMF^{2})$. Summing these terms gives
$O\!\left(BT\,(NF^{2}+MF^{2}+M^{2}F)\right)\approx O(BTM^{2}F),$ (30)
since $M^{2}F>MF^{2}>NF^{2}$.
Therefore, the overall complexity of the dual-view encoder combined with the high-order decoder amounts to $O(BTM^{2}F)$. Notably, for the downstream tasks, we only employ the dual-view encoder, reducing the practical complexity to $O(BTNF^{2})$.
Appendix I Interpretability analyses
Our method achieves the largest gains in Emotion, Gambling, Language, and Working Memory, indicating that the learned embedding is particularly effective at capturing higher-order cognitive processes. These tasks rely on distributed cross-network interactions, latent affective or value states, semantic integration, and sustained executive control. The consistent improvements suggest that our embedding encodes a more abstract and cognitively aligned representation—one that goes beyond local activation or static connectivity and instead reflects the dynamic, integrative computations underlying advanced human cognition.